el M Chapter 17 HW × tps Question 9 (of 20) 9 10.00 points Problem 17-20 Preferred stock dividends in arrears and valuing common stock [LO17-51 Storage Company has 590,000 shares of cumulative preferred stock outstanding, which has a stated dividend of $5.65. It is six years in amrears in its dividend payments Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods a. How much in total dollars is the company behind in its payments? (Do not round intermediate calculations. Input your answer in dollars, not millions (e.g. $1,234,000).) in b. The firm proposes The common stock will pay the following dividends over the next four years to offer new common stock to the preferred stockholders to wipe out the deficit О, $190 D 2.00 2.10 ENG 10 AM 1/12/2019 Type here to search
2.00 2.10 The is 10 percent. calculations and round your answer to 2 decimal places) and round your answer to the nearest whole share.) O Type here to search
Answer #1

a.Dividend per share = $5.65

Number of years = 6

Total Dividend in arrears = 590,000*5.65*6 = $20,001,000

b.Value of common stock = present value of dividends + present value of stock price

Stock price after 4 years = EPS after 4 years*Price Earning Ratio

= $4.24*13 = $55.12

= 1.90*PVF(10%, 1 year) +2.00*PVF(10%. 2 years) + 2.10*PVF(10%, 3years) + 2.20*PVF(10%, 4 years) + 55.12*PVF(10%, 4 years)

= $1.90*0.909 + 2.00*0.826 + 2.10*0.751 + 2.2*0.683 + 55.12*0.683

= $44.11

c. Number of shares required = $20,001,000/44.11

= 453,435 (approx.)